OpenStudy (anonymous):

What is the average rate of change for the function x F(x) 0 22.00 1 21.56 2 21.13

OpenStudy (anonymous):

Over which interval? \([0,1]\)? \([0,2]\)? \([1,2]\)?

OpenStudy (anonymous):

I think the entire thing

OpenStudy (anonymous):

Okay, given an interval \([a,b]\), the average rate of change of a function that is defined and continuous on that interval is given by \[\frac{f(b)-f(a)}{b-a}\] If we're using \([0,2]\), then the average r.o.c. will be \[\frac{F(2)-F(0)}{2-0}=\frac{21.13-22.00}{2}=-0.435\]